OpenStudy (anonymous):

Solve for the variable. -22 = x - 7

OpenStudy (anonymous):

x= -29

OpenStudy (anonymous):

all these questions you are asking are the same concept.. do it yourself after we have explained how to do it...

OpenStudy (anonymous):

no it is not -29

OpenStudy (anonymous):

-15

OpenStudy (ashleyisakitty):

the correct answer is -15. Try to do some on your own now! It's an easy concept once you get the hang of it :) Good luck!

OpenStudy (anonymous):

very easy

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

-22 = x - 7 -22 + 7 = x - 7 + 7 ... Add 7 to both sides. -15 = x + 0 -15 = x x = -15

OpenStudy (anonymous):

never mind its -15